The mathematical functions of supply and demand for broilers in the local market are given below:

The Supply function is:                 Qs = 2100 + 344P;

The Demand function is:              Qd = 3660 – 324P;

The Equilibrium Price is P.


a)     Determine the Equilibrium Price, P.



b)     Mathematically determine the equilibrium Supply and Demand volume for broilers in this market. Show all calculations.



1
Expert's answer
2022-06-08T17:24:57-0400

"Q_s=2100+344P,"

"Q_d=3660-324P,"

"Q_s=Q_d,"

"2100+344P=3660-324P,"

"668P=1560,"

"P=2.3,"

"Q_s=Q_d=2903.4."